# Unraveling Zeroth-Order Optimization through the Lens of Low-Dimensional Structured Perturbations

### v1

@@ 4.1 Influence of Effective Overlap on Convergence
4.2 Comparison of Low-Dimensional Constraints: Practical Implementations
Figure 2: Effective overlap $\rho$ for different methods under a block-diagonal Hessian with heterogeneous eigenspectra. Each boxen plot shows the distribution of effective overlap across 1000 trials for a given $\mathrm{srank}(M)$. The horizontal blue dashed lines indicate the mean effective overlap for each method. The same figure presented in log-scale and strip plot are provided in Appendix F.1.
We now explore three practical approaches to incorporate low-dimensional structure into the matrix $\mathbf{M}$ and examine their behavior. Specifically, we consider:
  1. 1.
    Low-rank projection: $\mathbf{M}=\mathbf{U}\mathbf{U}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{d\times s}$ is a random orthogonal matrix of rank $s<d$.
  2. 2.
    Sparse masking: $\mathbf{M}=\mathrm{diag}(\mathbf{m})$, where $\mathbf{m}=(m_{1},\ldots,m_{d})$ and $m_{i}\sim\mathrm{Bernoulli}(p),p\in(0,1)$.
  3. 3.
    Block sparse masking: $\mathbf{M}=\mathrm{diag}(\mathbf{m})$, but partition the parameters into $N$ blocks $\{B_{1},\ldots,B_{N}\}$. For a mask $\mathbf{m}=(m_{1},\cdots,m_{d})\in\mathbb{R}^{d}$, we set $m_{i}=1$ if $i\in B_{j}$ where $j$ is uniformly sampled at random from $[N]$ (This is reminiscent of block coordinate descent).
Problem Setup. To more closely mirror the behavior of neural network training, we adopt a block-diagonal Hessian with heterogeneous eigenspectra. Prior studies (Collobert, 2004; Zhang et al., 2024b, c) have shown, both theoretically and empirically, that neural network Hessians can be approximated by a block-diagonal matrix, with each dense block exhibiting a distinct eigenvalue distribution. Accordingly, we set the problem dimension to $d=1024$ and partition $\mathbf{H}$ into $16$ diagonal blocks (each of rank $16$). The eigenvalues in each block are set by randomly choosing a reference from $\{10,40,70,100\}$ and sampling each eigenvalue as an integer within ±$\pm 2$ of it.
To investigate a range of stable ranks $\mathrm{srank}(\mathbf{M})$, we systematically the hyperparameters of each method in a straightforward manner. Specifically, for the low-rank projection, we directly set the rank of $\mathbf{U}$; for the sparse masking, we choose $p=\mathrm{srank}(\mathbf{M})/d$; and for the block sparse masking, we select $N$ such that $d/N=\mathrm{srank}(\mathbf{M})$. Additional details on the experimental setup can be found in Appendix E.1.2.
@@ 4.2 Comparison of Low-Dimensional Constraints: Practical Implementations
However, the variance of $\rho$ differs substantially. In particular, block sparse masking exhibits a wider spread of effective overlaps. This behavior arises because the Hessian is block-diagonal, and each block’s eigenspectrum can differ considerably. Consequently, the effective overlap depends heavily on which block is selected at each iteration: choosing a block with large eigenvalues leads to a significantly higher $\rho$ than choosing one dominated by smaller eigenvalues.
This variance pattern underscores a critical insight: a higher variance translates into a greater opportunity to discover more effective substructures, particularly those with larger overlap. As a result, if substructure selection is guided by a heuristic or theoretical principle (rather than being purely random), block sparse masking can achieve a more pronounced performance boost. Moreover, block sparse masking significantly simplifies substructure selection. Concretely, low-rank projection requires identifying a handful of critical directions from uncountably many in the whole parameter space, and also sparse masking must consider an extensive set of coordinate combinations. In contrast, block sparse masking merely requires deciding which pre-defined block to activate, an approach that becomes increasingly advantageous as model size grows.
These two advantages, greater potential for improvement and simpler substructure selection, position block sparse masking as a promising approach for zeroth-order optimization in large-scale neural network training, where both ease of tuning and substantial performance gains are essential.
5 Experiments on Large Language Models
Table 1: Performance of fine-tuning OPT-13B on 1000 examples across various fine-tuning tasks. ICL denotes in-context learning, while FT represents full fine-tuning using first-order Adam (Kingma & Ba, 2015). The best results, excluding FT, are highlighted in bold. Results for methods other than MeZO-BCD are taken from Chen et al. (2024).
@@ 5 Experiments on Large Language Models
Task type ———————— classification ———————— – multiple choice – — generation —
@@
MeZO $91.3$ $68.2$ $66.1$ $68.1$ $61.5$ $59.4$ $59.4$ $88.0$ $81.3$ $81.8$ $31.3$
LoZO $91.7$ $70.4$ $69.6$ $71.9$ $63.5$ $60.8$ $63.0$ $\mathbf{89.0}$ $81.3$ $\mathbf{84.9}$ $30.7$
MeZO-BCD $\mathbf{93.0}$ $\mathbf{72.6}$ $\mathbf{71.4}$ $\mathbf{72.4}$ $\mathbf{64.4}$ $\mathbf{61.9}$ $\mathbf{63.2}$ $\mathbf{89.0}$ $\mathbf{81.7}$ $83.9$ $\mathbf{31.4}$
@@
In this section, we investigate the performance comparison of three approaches in LLM fine-tuning scenarios. Specifically, we compare LoZO (Chen et al., 2024) (employing low-rank projection), SparseMeZO (Liu et al., 2024b) (uses sparse masking), and our MeZO-BCD implementation, respectively. For MeZO-BCD, we treat a single transformer decoder layer (encompassing both self-attention and MLP components) as one block. At each optimization step, we either randomly choose the block to be updated or adopt a simple “flip-flop” strategy that alternates between ascending and descending block indices. Further details on MeZO-BCD is provided in Appendix D (refer to Algorithm 12).
Fine-tuning on OPT-13B. To evaluate how MeZO-BCD compares with existing methods [^1] Since SparseMeZO’s official implementation is unavailable, and our implementation cannot fully reproduce its reported performance in (Liu et al., 2024b), we exclude this baseline to maintain a fair comparison., we fine-tune OPT-13B (Zhang et al., 2022a) on multiple tasks, including SuperGLUE (Wang et al., 2019), SQuAD (Rajpurkar et al., 2018), and DROP (Dua et al., 2019). As shown in Table 1, MeZO-BCD surpasses both LoZO and MeZO under these configurations. Notably, this does not imply that MeZO-BCD categorically outperforms its counterparts when extensive hyperparameter tuning is performed on the latter. Rather, our results indicate MeZO-BCD to be at least on par with state-of-the-art ZO methods, suggesting it as a promising method for further exploration in large-scale optimization.
Figure 3: Comparison of training loss curves in terms of iterations (left) and wall-clock time (right) on OPT-1.3B fine-tuning with SST-2 for different methods.
Discussion on Efficiency. Figure 3 compares the convergence of fine-tuning OPT-1.3B (Zhang et al., 2022a) on SST-2 (Socher et al., 2013) in terms of both iterations (left) and wall-clock time (right). While LoZO, SparseMeZO, and MeZO-BCD converge at roughly similar rates per iteration, the differences in wall-clock time per iteration is significant. MeZO and LoZO exhibit similar wall-clock time with SparseMeZO being slightly slower. MeZO-BCD, in contrast is substantially faster by up to ×$\times\textbf{2.09}$. This benefit arises as MeZO-BCD only needs to sample random perturbations for a single block at each step.
From a memory perspective, MeZO-BCD does not require additional storage compared to MeZO, yielding similar memory benefits. In contrast, LoZO allocates extra memory for subspace maintenance across multiple steps (lazy subspace sampling), and SparseMeZO either stores sparse masks or incurs additional computation to reconstruct them each time. Hence, MeZO-BCD demonstrates superior efficiency in both run-time speed and memory usage.
Additionally, MeZO-BCD can offer unique advantages when combined with stateful optimizers such as Adam or momentum-based methods. Inspired by BAdam (Luo et al., 2024), one could optimize a selected block for a fixed number of steps while maintaining optimizer states only for that block, thereby significantly reducing memory overhead. Although we leave such exploration for future work, this possibility further highlights the efficiency advantages of MeZO-BCD in optimizing large foundation models.
6 Conclusion and Limitations
In this work, we introduced a unified theoretical framework for zeroth-order optimization that reveals how low-dimensional constraints, characterized by stable rank and effective overlap, reduce gradient noise, accelerate convergence, and enhance generalization. Empirical studies validated this framework and highlighted the potential of block coordinate descent, particularly due to its higher variance in effective overlap. Furthermore, we showed that block coordinate descent (MeZO-BCD) is a promising solution for large-scale ZO optimization, offering improved efficiency while achieving performance comparable to existing alternatives in fine-tuning LLM tasks. Despite these advances, several challenges remain as limitations: searching better block-selection strategies, exploring richer structural constraints, and incorporating adaptive gradients, which we leave as future work.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are may be potential societal consequences of our work. Specifically, with the growing use cases of the LLM fine-tuning, detailed insights on the ZO and potential efficient alternatives as well proposed in the paper, would help reduce the carbon footprint of such services.
@@ Appendix A Related Works
Theoretical analysis of zeroth-order (ZO) optimization has been an active research area, with extensive work on its convergence properties across various settings. Early studies established fundamental complexity results for both first-order and zeroth-order methods using Gaussian smoothing (Ghadimi & Lan, 2013). Later, optimal convergence bounds for zeroth-order stochastic gradient descent (SGD) were derived for convex objectives (Duchi et al., 2015) and extended to Gaussian-smoothing-based approaches (Nesterov & Spokoiny, 2017). In nonconvex settings, variance-reduced ZO optimization methods such as ZO-SVRG and ZO-SPIDER were introduced to improve convergence rates and function query efficiency (Ji et al., 2019). More recent works have analyzed ZO optimization under finite-difference, linear interpolation, and various smoothing schemes (Berahas et al., 2022), while others have investigated its behavior in non-smooth regimes, providing sample complexity bounds (Zhang et al., 2020; Davis et al., 2022) and demonstrating that non-smooth ZO optimization is not necessarily more difficult than the smooth case (Rando et al., 2023; Kornowski & Shamir, 2024).
@@
Zeroth-order (ZO) optimization, a long-standing technique in conventional machine learning (Spall, 1992; Ghadimi & Lan, 2013), was first explored for large language model (LLM) fine-tuning by MeZO (Malladi et al., 2023). By generating perturbations dynamically using random seeds, MeZO eliminates the need to store large perturbation matrices, significantly reducing memory overhead. This allows ZO optimization to operate with memory requirements comparable to inference. While MeZO provides theoretical guarantees, suggesting that scaling the learning rate with respect to the problem dimension ensures dimension-free convergence, this approach inherently results in slow convergence.
Several approaches have been proposed to improve the efficiency of ZO optimization. MeZO-SVRG (Gautam et al., 2024) integrates the SVRG algorithm to reduce gradient variance, leading to faster convergence and improved fine-tuning performance. SparseMeZO (Liu et al., 2024b) enhances computational efficiency by updating only a subset of parameters, while Fisher-informed sparsity (Guo et al., 2024) prioritizes updates based on Fisher information. A benchmarking study (Zhang et al., 2024d) evaluates various ZO methods, including SGD, SignSGD, and Adam. More recently, LOZO (Chen et al., 2024) introduced low-rank perturbations, leveraging the low intrinsic dimensionality of gradients in LLMs to improve convergence and fine-tuning performance.
@@
Appendix B Proof of Convergence
In this section, we provide the convergence analysis in details. To this end, we revisit our zeroth-order gradient estimator and the update rule.
@@ B.1 Preliminary
Note that $\mathcal{L}_{\mu}$ approaches to the original loss $\mathcal{L}$ as $\mu\rightarrow 0$. This indicates that the estimation error between $\mathcal{L}_{\mu}$ and $\mathcal{L}$ or the distance of gradients between $\nabla\mathcal{L}_{\mu}$ and $\nabla\mathcal{L}$ would depend on the smoothing parameter $\mu$.
@@
for simplicity. For the positive semi-definite matrix $\mathbf{M}_{t}$ with $\mathrm{srank}(\mathbf{M}_{t})\leq s\ll d$, we have
@@
Under the maximal difference $\mu L\sqrt{2(d+s)}$ between $\nabla\mathcal{L}_{\mu,\mathbf{M}_{t}}(\bm{\theta})$ and $\nabla\mathcal{L}_{\mu}(\bm{\theta})$, the inner product could be easily lower-bounded by
@@
B.2 Technical Lemmas for Matrix Calculus
@@
Lemma B.3 (Nesterov & Spokoiny (2017)).
@@
$=\left(\sum\limits_{k=1}^{d}\Sigma_{kk}\right)\left(\sum\limits_{i,j}a_{i}a_{j}\Sigma_{ij}\right)$
$=\mathrm{Tr}(\bm{\Sigma})\mathbf{a}^{\mathsf{T}}\bm{\Sigma}\mathbf{a}$
$=\mathrm{Tr}\big{(}\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\big{)}\mathbf{a}^{\mathsf{T}}\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{a}$
$=\mathrm{Tr}\big{(}\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\big{)}\big{\|}\mathbf{Q}^{\mathsf{T}}\mathbf{a}\big{\|}^{2}$
@@
$=2\sum\limits_{k=1}^{d}\sum\limits_{i,j}a_{i}\Sigma_{ik}a_{j}\Sigma_{jk}$
$=2\sum\limits_{k=1}^{d}\left(\sum\limits_{i=1}^{d}a_{i}\Sigma_{ik}\right)\left(\sum\limits_{j=1}^{d}a_{j}\Sigma_{jk}\right)$
$=2\sum\limits_{k=1}^{d}[\bm{\Sigma}\mathbf{a}]_{k}^{2}$
$=2\|\bm{\Sigma}\mathbf{a}\|^{2}$
$=2\mathbf{a}^{\mathsf{T}}\bm{\Sigma}^{2}\mathbf{a}$
@@
B.3 Technical Lemmas for Gaussian Smoothed Loss

### v2

@@ 4.3 Experiments on LLM fine-tuning
[…]
MeZO-BCD $\mathbf{0.0339~{}(\times 1.83)}$ $\mathbf{0.0637~{}(\times 2.10)}$ $\mathbf{0.375~{}(\times 2.77)}$ $25.3$ $33.6$ $43.8$
Figure 2: Comparison of training loss curves in terms of iterations (left) and wall-clock time (middle) on OPT-1.3B (Zhang et al., 2022a) fine-tuning with SST-2 (Socher et al., 2013) for different methods.
To empirically demonstrate the practical advantages of MeZO-BCD, particularly its enhanced efficiency and competitive performance, we conduct extensive fine-tuning experiments on LLMs. Key results are presented in Figure 2 and Table 12. Detailed experimental setups, including datasets, hyperparameters and model specifics are provided in Appendix E.2.
MeZO-BCD exhibits superior wall-clock time efficiency.  The most significant empirical validation of our approach is MeZO-BCD’s superior wall-clock time efficiency, as shown in Figure 2 (Right) and Table 1. Across different model scales, MeZO-BCD consistently has less iteration time compared to other ZO methods like MeZO, LoZO, and SparseMeZO. Notably, as the model size increases, the acceleration gain becomes higher due to the increased parameter loading overhead. This directly substantiates our central argument from Section 4.1 that MeZO-BCD’s design significantly reduces memory loading overhead, translating into tangible training speedups.
MeZO-BCD incurs no additional memory compared to MeZO.  MeZO-BCD operates under the same forward-only regime as MeZO but perturbs and updates only a single block at each step. As a result, it avoids any increase in memory footprint. This is empirically confirmed in Table 1, where the peak memory remains effectively identical across different methods [^1] Minor differences in peak memory across methods (e.g., on SST-2) are due to system-level factors such as CUDA memory fragmentation or transient buffer allocation, rather than algorithmic differences..
Table 2: Performance of fine-tuning OPT-13B on 1000 examples across various tasks. ICL denotes in-context learning, while FT denotes full fine-tuning via first-order Adam (Kingma and Ba, 2015). The best results except FT are highlighted in bold. Results for baselines are taken from LoZO (Chen et al., 2024).
@@
Task type —————————– classification —————————– – multiple choice – — generation —
@@
MeZO $91.3$ $68.2$ $66.1$ $68.1$ $61.5$ $59.4$ $59.4$ $88.0$ $81.3$ $81.8$ $\mathbf{31.3}$
LoZO $91.7$ $\mathbf{70.4}$ $69.6$ $71.9$ $\mathbf{63.5}$ $60.8$ $\mathbf{63.0}$ $\mathbf{89.0}$ $81.3$ $\mathbf{84.9}$ $30.7$
MeZO-BCD $\mathbf{93.0}$ $70.0$ $\mathbf{71.4}$ $\mathbf{72.4}$ $\mathbf{63.5}$ $\mathbf{61.9}$ $62.8$ $\mathbf{89.0}$ $\mathbf{81.6}$ $83.7$ $31.0$
@@
MeZO-BCD shows competitive convergence and fine-tuning performance.  Notably, this efficiency gain does not compromise iteration complexity or fine-tuning performance. As shown by the training loss curves in Figure 2 (left), MeZO-BCD converges on par with baseline ZO methods. Furthermore, Table 2 demonstrates MeZO-BCD’s strong fine-tuning performance on a diverse suite of downstream tasks with an OPT-13B (Zhang et al., 2022a) model, where it frequently outperforms other ZO methods (Malladi et al., 2023; Chen et al., 2024) [^2] Since SparseMeZO’s official implementation is unavailable, and our implementation cannot fully reproduce its reported performance in (Liu et al., 2024b), we exclude this baseline to maintain a fair comparison.. Additional experimental results are provided in Appendix F.2.
5 Discussion
5.1 MeZO-BCD as a Scalable Foundation for Zeroth-Order Optimization
Adaptive Block Selection.
While our theoretical analysis (Propositions 3.4 and 3.6) shows that various subspace perturbation schemes share similar expected convergence rates, their subspace alignment $\rho$ can differ substantially in distribution. This highlights a natural opportunity for improvement via adaptive selection of subspaces (specifically, blocks in the case of MeZO-BCD). Compared to low-rank or sparse perturbation, where adaptivity requires searching over continuous subspaces or combinatorial coordinate subsets, MeZO-BCD operates on a small set of architecturally defined, semantically meaningful blocks. This makes adaptive scheduling more tractable and interpretable in practice. Moreover, since blocks often differ in curvature or gradient scale across layers in LLMs (Zhang et al., 2024b, c), MeZO-BCD provides a well-suited interface for learning dynamics-aware scheduling. We leave the empirical realization of this idea as promising future work.
Integration with Stateful Optimizers.
Stateful optimizers such as Adam (Kingma and Ba, 2015) have been instrumental in first-order optimization, but their extension to zeroth-order methods has shown limited benefits. Previous works (Malladi et al., 2023; Zhang et al., 2024e) found marginal performance improvements and substantial memory overhead due to per-parameter optimizer state storage. MeZO-BCD offers a new path forward: by perturbing and updating only one block at a time, optimizer states can be stored and updated locally for the active block, without incurring full-model memory costs. This enables practical and scalable use of stateful optimizers in ZO settings, potentially opening a new line of research into adaptive gradient-free methods that leverage momentum, variance correction, or block-wise learning rates in a resource-efficient way. Preliminary investigations on the future directions are provided in Appendix G.
5.2 Limitations
Despite its scalability and extensibility, MeZO-BCD has several limitations. First, while our theoretical framework provides general insights into the role of subspace perturbations and dimensionality, it does not offer task-specific guidance on optimal perturbation structures or block sizes. Second, although we highlight promising extensions including adaptive block selection and integration with stateful optimizers, this work focuses on establishing the core algorithm and validating its foundational benefits. Empirical investigation of these directions remains an important avenue for future work.
6 Conclusion
In this work, we introduced a unified theoretical framework for understanding zeroth-order optimization under subspace perturbations, connecting convergence and generalization through the notion of subspace alignment. Our analysis provides new insights into how different perturbation strategies behave and what distinguishes them beyond that. Building on the theoretical insights, we proposed MeZO-BCD, a block coordinate descent method for zeroth-order optimization that achieves substantial practical speedups. Our experiments on LLM fine-tuning confirm its efficiency and competitiveness, and its structured design opens promising directions for future research, including adaptive block selection and integration with stateful optimizers.
@@ 6 Conclusion
[…]
Broader Impact Statement
This work improves the efficiency and scalability of zeroth-order optimization for fine-tuning large language models, offering a practical alternative to gradient-based approaches in resource-constrained settings. By lowering computational and memory demands, it may help reduce the environmental footprint of large-scale AI training and broaden access to advanced machine learning technologies. However, expanding access to powerful fine-tuning tools also raises societal concerns, including the risk of misuse or unintended consequences from customized language models. We advocate for responsible usage, with attention to transparency, fairness, and ongoing evaluation of social impact as such methods are deployed in practice.
@@ Appendix A Related Works
Theoretical analysis of zeroth-order (ZO) optimization has been an active research area, with extensive work on its convergence properties across various settings. Early studies established fundamental complexity results for both first-order and zeroth-order methods using Gaussian smoothing (Ghadimi and Lan, 2013). Later, optimal convergence bounds for zeroth-order stochastic gradient descent (SGD) were derived for convex objectives (Duchi et al., 2015) and extended to Gaussian-smoothing-based approaches (Nesterov and Spokoiny, 2017). In nonconvex settings, variance-reduced ZO optimization methods such as ZO-SVRG and ZO-SPIDER were introduced to improve convergence rates and function query efficiency (Ji et al., 2019). More recent works have analyzed ZO optimization under finite-difference, linear interpolation, and various smoothing schemes (Berahas et al., 2022), while others have investigated its behavior in non-smooth regimes, providing sample complexity bounds (Zhang et al., 2020; Davis et al., 2022) and demonstrating that non-smooth ZO optimization is not necessarily more difficult than the smooth case (Rando et al., 2023a; Kornowski and Shamir, 2024).
@@
Zeroth-order (ZO) optimization, a long-standing technique in conventional machine learning (Spall, 1992; Ghadimi and Lan, 2013), was first explored for large language model (LLM) fine-tuning by MeZO (Malladi et al., 2023). By generating perturbations dynamically using random seeds, MeZO eliminates the need to store large perturbation matrices, significantly reducing memory overhead. This allows ZO optimization to operate with memory requirements comparable to inference. While MeZO provides theoretical guarantees, suggesting that scaling the learning rate with respect to the problem dimension ensures dimension-free convergence, this approach inherently results in slow convergence.
Several approaches have been proposed to improve the efficiency of ZO optimization. MeZO-SVRG (Gautam et al., 2024) integrates the SVRG algorithm to reduce gradient variance, leading to faster convergence and improved fine-tuning performance. SparseMeZO (Liu et al., 2024b) enhances computational efficiency by updating only a subset of parameters, while Fisher-informed sparsity (Guo et al., 2024) prioritizes updates based on Fisher information. A benchmarking study (Zhang et al., 2024e) evaluates various ZO methods, including SGD, SignSGD, and Adam. More recently, LOZO (Chen et al., 2024) introduced low-rank perturbations, leveraging the low intrinsic dimensionality of gradients in LLMs to improve convergence and fine-tuning performance.
@@
Appendix B Proof for Convergence
In this section, we provide the theoretical analysis of Section 3 in detail.
B.1 Discussion on the Use of Assumption 3.1
Assumption 3.1 is well-justified in the context of large-scale language model optimization. A closely related, but even stronger, assumption was already adopted in MeZO (Malladi et al., 2023), where it was supported both theoretically and empirically. Specifically, MeZO justified this condition by noting that the Hessian of the loss for deep neural networks (especially during fine-tuning) often exhibits a remarkably low local effective rank (Papyan, 2018, 2020; Ghorbani et al., 2019; Wu et al., 2020; Sagun et al., 2017; Zhang et al., 2024c, b), and that even for billion-parameter models, zeroth-order optimization remains effective, as also evidenced by the rapid convergence observed in practice. These empirical phenomena are further supported by numerous studies demonstrating that the Hessian spectrum is highly concentrated near zero, with only a handful of large outliers providing an effective upper bound on the local rank.
Our formulation is strictly less restrictive than MeZO’s, as it only requires the assumption to hold within a local $\mathcal{O}(s)$-ball and allows for a strictly positive smoothing parameter $\mu$. Given that the original assumption was already validated in prior work, adopting this weaker version in our analysis is not only natural but also better aligned with the practical realities of large-scale model fine-tuning.
B.2 Preliminaries
Recall that our zeroth-order gradient estimator and the subspace ZO-SGD.
@@ B.2 Preliminaries
Note that it is clear that $\mathcal{L}_{\mu}$ approaches to the original loss $\mathcal{L}$ as $\mu\rightarrow 0$. This indicates that the estimation error between $\mathcal{L}_{\mu}$ and $\mathcal{L}$ or the distance of gradients between $\nabla\mathcal{L}_{\mu}$ and $\nabla\mathcal{L}$ would depend on the smoothing parameter $\mu$, thus the key point is “how to handle this mismatch”.
@@
for simplicity. For the positive semi-definite matrix $\mathbf{M}_{t}$ with $\mathrm{srank}(\mathbf{M}_{t})\leq s$, it follows that
@@
Under the maximal difference $\mu L\sqrt{2(d+s)}$ between $\nabla\mathcal{L}_{\mu,\mathbf{M}_{t}}(\bm{\theta})$ and $\nabla\mathcal{L}_{\mu}(\bm{\theta})$, it is easy to derive that the inner product is lower-bounded as
@@
B.3 Technical Lemmas for Matrix Calculus
@@
Lemma B.3 (Lemma 1 in (Nesterov and Spokoiny, 2017)).
@@
$=\left(\sum\limits_{k=1}^{d}\Sigma_{kk}\right)\left(\sum\limits_{i,j}a_{i}a_{j}\Sigma_{ij}\right)=\mathrm{Tr}(\bm{\Sigma})\mathbf{a}^{\mathsf{T}}\bm{\Sigma}\mathbf{a}$
$=\mathrm{Tr}\big{(}\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\big{)}\mathbf{a}^{\mathsf{T}}\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{a}=\mathrm{Tr}\big{(}\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\big{)}\big{\|}\mathbf{Q}^{\mathsf{T}}\mathbf{a}\big{\|}^{2}$
@@
$=2\sum\limits_{k=1}^{d}\sum\limits_{i,j}a_{i}\Sigma_{ik}a_{j}\Sigma_{jk}=2\sum\limits_{k=1}^{d}\left(\sum\limits_{i=1}^{d}a_{i}\Sigma_{ik}\right)\left(\sum\limits_{j=1}^{d}a_{j}\Sigma_{jk}\right)$
$=2\sum\limits_{k=1}^{d}[\bm{\Sigma}\mathbf{a}]_{k}^{2}=2\|\bm{\Sigma}\mathbf{a}\|^{2}=2\mathbf{a}^{\mathsf{T}}\bm{\Sigma}^{2}\mathbf{a}$
@@
B.4 Technical Lemmas for Gaussian Smoothed Loss